OpenStudy (anonymous):

Find a. A.6 b.6 times the square root of 2 c. 3 time the square root of 2 d. 3

OpenStudy (anonymous):

is it b or c?

OpenStudy (saifoo.khan):

Use sin ratio.\[\sin \theta = \frac{opposite}{hypotenuse} \]

OpenStudy (anonymous):

you will get an approximate answer using @saifoo.khan method so you have to carry out the multiplication in b and c to see which one it is.